As a financial analyst for Wavigations, you are tasked with raising equity capital. The CEO would like to know by how much the cost of new stock would exceed the cost of common from reinvested earnings (in percent). Wavigations’s common stock currently sells for $78 per share,the company expects to earn $6 per share during the current year, its expected payout ratio is 45%, and its expected constant growth rate is 3%. New stock can be sold to the public at the current price, but a flotation cost of 4% would be incurred.
Answer #1

Cost of new stock-Cost of reinvested earnings=Expected Earnings*Payout Ratio/Current Price*(1/(1-flotation cost)-1)=6*45%/78*(1/(1-4%)-1)=0.144%